# Unsupervised Domain Adaptation for Binary Classification with an Unobservable Source Subpopulation

### v2

@@ Abstract
We study an unsupervised domain adaptation problem where the source domain consists of subpopulations defined by the binary label $Y$ and a binary background (or environment) $A$. We focus on a challenging setting in which one such subpopulation in the source domain is unobservable. Naively ignoring this unobserved group can result in biased estimates and degraded predictive performance. Despite this structured missingness, we show that the prediction in the target domain can still be recovered. Specifically, we rigorously derive both background-specific and overall prediction models for the target domain. For practical implementation, we propose the distribution matching method to estimate the subpopulation proportions. We provide theoretical guarantees for the asymptotic behavior of our estimator, and establish an upper bound on the prediction error. Experiments on both synthetic and real-world datasets show that our method outperforms the naive benchmark that does not account for this unobservable source subpopulation.
Unsupervised Domain Adaptation for Binary Classification with an Unobservable Source Subpopulation
@@ 1 Introduction
In this work, we focus on a more challenging and practically relevant UDA setting where a structured subpopulation is entirely missing from the source domain. Specifically, we consider binary label $Y$ and a binary background or environment variable $A$, and study the case where one subpopulation, defined by a particular combination of $Y$ and $A$, is unobserved in the source. This structured missingness is not merely a sampling artifact, but often reflects real-world constraints in data collection. For instance, in the widely studied Waterbirds dataset (Sagawa et al., 2019), waterbirds ($Y=1$) photographed in water environments ($A=1$) can be rare or entirely absent due to the difficulty of capturing such images in the wild. This issue arises in many other disciplines as well. In healthcare, certain patient subgroups, defined jointly by disease status and demographics, may be underrepresented or absent in historical datasets due to restrictive inclusion criteria or changes in clinical practice over time. When such models are applied to broader populations, unobserved subgroups can suffer from systematic mispredictions. This structured missingness (Mitra et al., 2023) fundamentally changes some statistical properties when comparing the source and target domains, and, if unaddressed, can lead to severely biased estimation and unreliable prediction in the target domain. These structured gaps pose new challenges that are not adequately addressed by conventional UDA techniques, which motivates our work.
@@
We validate our approach through experiments on both synthetic and real-world datasets. We simulate domain adaptation scenarios where one subpopulation is systematically excluded from the source data and evaluate our method against baseline approaches that do not account for this missing group. Across a range of settings, our method consistently achieves higher accuracy and F1 scores, particularly on the subpopulation absent from the source. These results highlight the practical value of explicitly modeling structured missingness and demonstrate that our approach leads to more reliable predictions in the target domain. To summarize, this paper makes the following novel contributions:
@@ 2 Related Work
OOD generalization refers to the ability of a prediction model to perform well on test data drawn from a distribution that differs from the training data. In our context, the subpopulation $(Y=1,A=1)$ in the target can be regarded as the OOD data while the other three subpopulations are in-distribution data. For a comprehensive overview of OOD generalization, we refer the readers to the excellent survey (Liu et al., 2021), which reviewed real-world datasets, evaluation protocols, and key challenges in this area. In the OOD generalization literature, different methods were proposed with different emphases: (Arjovsky et al., 2019) emphasized the need to minimize invariant risk across different environments to ensure consistent model performance, whereas (Sagawa et al., 2019) underscored the importance of distributionally robust optimization (DRO) and various regularization techniques in reducing performance disparities across subgroups. In addition, (Bahng et al., 2020) introduced adversarial training as a method for learning de-biased representations, which is critical for promoting fairness in machine learning models, and (Sohoni et al., 2020) examined the issue of robustness in classification tasks involving coarse classes that contain finer subclasses, enhancing model performance across all subclasses.
@@
OOD detection is the task of identifying inputs at test time that do not come from the same distribution as the training data. Its goal is to prevent a model from making confident but incorrect predictions on unfamiliar or anomalous inputs by flagging them as OOD. There are a variety of techniques developed for OOD detection in the literature. For example, (Hendrycks and Gimpel, 2017) introduced a simple yet effective method for detecting both misclassified and OOD inputs in neural networks. Liang et al. (2018) (ODIN) proposed an improved method for detecting OOD inputs by applying temperature scaling to the softmax outputs and adding small input perturbations during inference. ODIN significantly outperformed previous baseline methods, including the maximum softmax probability approach, and set a new standard for OOD detection in classification tasks. Other techniques include but not limited to, outlier exposure (Hendrycks et al., 2018; Papadopoulos et al., 2021), ConfGAN (Sricharan and Srivastava, 2018) and OodGAN (Marek et al., 2021). In addition, Fort et al. (2021) provided an extensive empirical study of OOD detection methods across a wide range of datasets, architectures, and training regimes.
@@
Spurious correlation is a major obstacle to OOD generalization, where models often rely on non-causal features that can degrade performance, particularly when these correlations do not generalize across domains. For example, a model trained to classify cows might rely on green pastures (background) instead of the cow itself. On a desert background, it fails. This is also the case in the Waterbirds dataset where the spurious correlation exists between label $Y$ and background $A$. Different learning strategies were proposed to discover and mitigate the impact of spurious correlation on model performance, as well as to improve model robustness. For example, (Wu et al., 2023) introduced an attention-based approach to automatically identify spurious concepts and apply adversarial training to reduce reliance on them. Another approach proposed by (Kumar et al., 2023) used causal regularization to detect and discourage spurious dependencies, allowing for scalable robustness across shifts. In addition, (Sagawa et al., 2020) investigated why overparameterization exacerbates spurious correlations, and (Kirichenko et al., 2022) found that retraining only the final layer on a small, balanced dataset can restore robustness against spurious correlations. Also, (Wang and Wang, 2024) developed a theoretical model to analyze the influence of spurious correlation strength, sample size, and feature noise on learning. Spurious correlations were also investigated in feature learning (Izmailov et al., 2022; Qiu et al., 2024), reinforcement learning (Ding et al., 2023), OOD detection (Ming et al., 2022), and text classification (Wang and Culotta, 2020). One can also resort to a comprehensive survey paper (Ye et al., 2024) on this topic.
@@ 3 Problem Setup and Notation
$R$ $Y$ AA $\bf X$} Sample Size Proportion Prediction Models
Source 1 1 1 0 1 0 1 0 0 $n_{101}$ $n_{110}$ $n_{100}$ $p_{101}=\alpha_{01}\pi$ $p_{110}=\alpha_{10}\pi$ $p_{100}=\alpha_{00}\pi$ $\xi_{1}({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},A=1,R=1)$
@@
Note that this assumption is made without loss of generality, as alternative combinations, such as $(Y=0,A=1)$, $(Y=1,A=0)$, or $(Y=0,A=0)$, can be similarly assumed to have zero probability. To characterize the distributional connection between the two domains, we impose a structured conditional invariance assumption:
@@
that is, the conditional distribution of features $\bf X$} given $(Y,A)$ remains the same across domains. This can be regarded as a conditional version, or, more nuanced version, of label shift where the marginal distribution of labels (now, the combination of both label and background) varies across domains (e.g., Du Plessis and Sugiyama, 2014; Garg et al., 2020; Iyer et al., 2014; Lipton et al., 2018; Nguyen et al., 2016; Tasche, 2017; Tian et al., 2023; Zhang et al., 2013). It indicates, conditional on background $A$, the label shift assumption holds. It is equivalent to $p(R|{\bf X},Y,A)=p(R|Y,A)$, the independence between $R$ and $\bf X$}, conditional on $(Y,A)$. In practice, this assumption may be suitable in many applications. Below we give two examples to illustrate the rationality of this assumption. For instance, we aim to predict user clicks on advertisements for a new batch of users (target domain, $R=0$) using historical data (source domain, $R=1$). Conditional on the advertisement type $A$ and whether the user clicks $Y$, the distribution of browsing behavior features $\bf X$} is assumed to remain stable across time periods. This is because user clicks are fundamentally determined by ad content and user interests, not by the time period in which data are collected. As another example, suppose we have datasets from two hospitals ($R=1$ indicates the source hospital and $R=0$ indicates the target hospital). Here, $\bf X$} represents imaging features, $Y$ is the disease type, and $A$ denotes patient attributes such as gender or age group. Then, conditional on the disease type $Y$ and demographic attributes $A$, the distribution of imaging features $\bf X$} is expected to remain the same across hospitals. This is because imaging characteristics for a given disease and demographic group are not systematically altered by the hospital. The main difference between hospitals lies in sampling proportions rather than in conditional distributions.
@@ 4 Proposed Methodology
Our goal in this work is to correctly identify and successfully implement, under our UDA setting, the two background-specific prediction models $\eta_{1}({\bf x})$ and $\eta_{0}({\bf x})$ and the overall prediction model $\eta({\bf x})$, in the target domain. All of the three models were precisely defined in Table 1.
4.1 The naive benchmark
As the benchmark, one may naively apply the three source domain prediction models $\xi_{1}({\bf x})$, $\xi_{0}({\bf x})$, and $\xi({\bf x})$ to the target. First, one can use the observed data to implement the overall source domain prediction model $\xi({\bf x})$ and one background-specific prediction model $\xi_{0}({\bf x})$ as
@@ 4.1 The naive benchmark
Afterwards, even though the subpopulation $(Y=1,A=1)$ is entirely absent in the source, still one can compute the other background-specific prediction model
$\xi_{1}({\bf x})$ $=\frac{\xi({\bf x})-\xi_{0}({\bf x})\{1-\tau_{1}({\bf x})\}}{\tau_{1}({\bf x})},\mbox{ where }$ (5)
$\tau_{1}({\bf x})$ $=\hbox{pr}(A=1\mid{\bf X}={\bf x},R=1),$ (6)
can also be implemented using the observed data.
@@ 4.2 Model adaptation from source to target
The most challenging aspect of this work is to adapt the model for the $A=1$ background since the component $(Y=1,A=1)$ is entirely absent in the source. Nevertheless, we can still correctly derive the three prediction models for the target domain, as shown below.
@@
Define the model $\tau_{0}({\bf x})=\hbox{pr}(A=1\mid{\bf X}={\bf x},R=0)$ and the model
@@
both of which can be implemented using the observed data in our UDA setting. Then the three prediction models in the target domain are given by:
@@
The proof of this result is provided in Appendix S.1. Proposition 1 illustrated that, in general, the naive method presented in Section 4.1 fails. There are no explicit relations between $\eta_{1}({\bf x})$ and $\xi_{1}({\bf x})$ or between $\eta({\bf x})$ and $\xi({\bf x})$. For the relation between $\eta_{0}({\bf x})$ and $\xi_{0}({\bf x})$, they coincide only in the special case that $\beta_{10}/\alpha_{10}=\beta_{00}/\alpha_{00}$, which corresponds to a proportionality condition between the class-conditional densities across domains. Outside of this narrow scenario, the naive approach systematically misestimates the target posterior, leading to biased predictions.
This result also implies that model adaptation fundamentally relies on estimating the proportions of key subgroups in the target population. In particular, for individuals with $A=1$, one only needs to estimate $\beta_{01}$, while for those with $A=0$, it suffices to estimate the ratio $\beta_{10}/\beta_{00}$. Denote ${\boldsymbol{\beta}}=(\beta_{10},\beta_{00})^{\rm T}$. It can be seen that, accurate estimation of the parameter $\boldsymbol{\beta}$} in the target domain enables valid model adaptation across domains. Before developing methods for estimating $\boldsymbol{\beta}$} in Section 4.4, we first present some model identification considerations.

### v3

@@ Abstract
We study an unsupervised domain adaptation problem where the source domain consists of subpopulations defined by the binary label $Y$ and a binary background (or environment) $A$. We focus on a challenging setting in which one such subpopulation in the source domain is unobservable. Naively ignoring this unobserved group can result in biased estimates and degraded predictive performance. Despite this structured missingness, we show that the prediction in the target domain can still be recovered. Specifically, we rigorously derive both background-specific and overall predictive probabilities for the target domain. For practical implementation, we propose the distribution matching method to estimate the subpopulation proportions. We provide theoretical guarantees for the asymptotic behavior of our estimator, and establish an upper bound on the prediction error. Experiments on both synthetic and real-world datasets show that our method outperforms the naive benchmarks that do not account for this unobservable source subpopulation properly.
@@ 1 Introduction
In many applications, the goal is to build a target-domain prediction model to predict a binary label $Y$ that identifies a specific object in an image (e.g., classifying a waterbird versus a landbird in the Waterbirds dataset). Additionally, the image typically contains contextual attributes represented by a binary background or environment variable $A$, such as whether the background consists of water or land. In this work, we focus on a more challenging and practically relevant UDA setting where a structured subpopulation is entirely missing from the source domain. Specifically, we study the case where one subpopulation, defined by a particular combination of $Y$ and $A$, is unobserved in the source. This structured missingness is not merely a sampling artifact, but often reflects real-world constraints in data collection. For instance, in the widely studied Waterbirds dataset (Sagawa et al., 2019), waterbirds ($Y=1$) photographed in water environments ($A=1$) can be rare or entirely absent due to the difficulty of capturing such images in the wild. This issue arises in many other disciplines as well. In healthcare, certain patient subgroups, defined jointly by disease status and demographics, may be underrepresented or absent in historical datasets due to restrictive inclusion criteria or changes in clinical practice over time. When such models are applied to broader populations, unobserved subgroups can suffer from systematic mispredictions. This structured missingness (Mitra et al., 2023) fundamentally changes some statistical properties when comparing the source and target domains, and, if unaddressed, can lead to severely biased estimation and unreliable prediction in the target domain. These structured gaps pose new challenges that are not adequately addressed by conventional UDA techniques, which motivates our work.
@@
We validate our approach through experiments on both synthetic and real-world datasets. We simulate domain adaptation scenarios where one subpopulation is systematically excluded from the source data and evaluate our method against baseline approaches that do not account for this missing group properly. Across a range of settings, our method consistently achieves higher accuracy and F1 scores. These results highlight the practical value of explicitly modeling structured missingness and demonstrate that our approach leads to more reliable predictions in the target domain. To summarize, this paper makes the following novel contributions:
@@ 2 Related Work
OOD generalization refers to the ability of a prediction model to perform well on test data drawn from a distribution that differs from the training data. In our context, the subpopulation $(Y=1,A=1)$ in the target can be regarded as the OOD data while the other three subpopulations are in-distribution data. For a comprehensive overview of OOD generalization, we refer the readers to the excellent survey (Liu et al., 2021), which reviewed real-world datasets, evaluation protocols, and key challenges in this area. In the OOD generalization literature, different methods were proposed with different emphases: Arjovsky et al. (2019) emphasized the need to minimize invariant risk across different environments to ensure consistent model performance, whereas Sagawa et al. (2019) underscored the importance of distributionally robust optimization (DRO) and various regularization techniques in reducing performance disparities across subgroups. In addition, Bahng et al. (2020) introduced adversarial training as a method for learning de-biased representations, which is critical for promoting fairness in machine learning models, and Sohoni et al. (2020) examined the issue of robustness in classification tasks involving coarse classes that contain finer subclasses, enhancing model performance across all subclasses.
@@
OOD detection is the task of identifying inputs at test time that do not come from the same distribution as the training data. Its goal is to prevent a model from making confident but incorrect predictions on unfamiliar or anomalous inputs by flagging them as OOD. There are a variety of techniques developed for OOD detection in the literature. For example, Hendrycks and Gimpel (2017) introduced a simple yet effective method for detecting both misclassified and OOD inputs in neural networks. Liang et al. (2018) (ODIN) proposed an improved method for detecting OOD inputs by applying temperature scaling to the softmax outputs and adding small input perturbations during inference. ODIN significantly outperformed previous baseline methods, including the maximum softmax probability approach, and set a new standard for OOD detection in classification tasks. Other techniques include but not limited to, outlier exposure (Hendrycks et al., 2018; Papadopoulos et al., 2021), ConfGAN (Sricharan and Srivastava, 2018) and OodGAN (Marek et al., 2021). In addition, Fort et al. (2021) provided an extensive empirical study of OOD detection methods across a wide range of datasets, architectures, and training regimes.
@@
Spurious correlation is a major obstacle to OOD generalization, where models often rely on non-causal features that can degrade performance, particularly when these correlations do not generalize across domains. For example, a model trained to classify cows might rely on green pastures (background) instead of the cow itself. On a desert background, it fails. This is also the case in the Waterbirds dataset where the spurious correlation exists between label $Y$ and background $A$. Different learning strategies were proposed to discover and mitigate the impact of spurious correlation on model performance, as well as to improve model robustness. For example, Wu et al. (2023) introduced an attention-based approach to automatically identify spurious concepts and apply adversarial training to reduce reliance on them. Another approach proposed by Kumar et al. (2023) used causal regularization to detect and discourage spurious dependencies, allowing for scalable robustness across shifts. In addition, Sagawa et al. (2020) investigated why overparameterization exacerbates spurious correlations, and Kirichenko et al. (2022) found that retraining only the final layer on a small, balanced dataset can restore robustness against spurious correlations. Also, Wang and Wang (2024) developed a theoretical model to analyze the influence of spurious correlation strength, sample size, and feature noise on learning. Spurious correlations were also investigated in feature learning (Izmailov et al., 2022; Qiu et al., 2024), reinforcement learning (Ding et al., 2023), OOD detection (Ming et al., 2022), and text classification (Wang and Culotta, 2020). One can also resort to a comprehensive survey paper (Ye et al., 2024) on this topic.
Imbalanced classification and few/zero-shot learning
In imbalanced classificaton, all classes are observed in the training data but appear with highly unequal frequencies, and many methods focus on reweighting or resampling strategies to improve performance, particularly on minority classes. Few-shot learning (Wang et al., 2020), such as one-shot learning (Li et al., 2006), refers to a learning paradigm in which a model learns from a very small number of labeled examples and then generalizes to novel classes. Unlike traditional supervised learning that requires large amounts of data, few-shot learning leverages a limited number of examples together with task-specific prior knowledge or structural information. Zero-shot learning (Wang et al., 2019) addresses tasks in which no labeled examples for the target classes are available during training, typically relying on auxiliary information such as semantic attributes or textual descriptions to relate seen and unseen classes. In contrast to these settings, our work considers a different challenge: a structured subgroup defined by a specific combination of $(Y,A)$ is completely absent from the source domain. This missing-subgroup setting introduces a distinct type of distribution shift that cannot be addressed directly by existing methods designed for imbalanced classification or few/zero-shot learning.
Adversatial domain adaptation
Adversarial domain adaptation methods, such as DANN (Ganin et al., 2016) and ADDA (Tzeng et al., 2017), aim to align the source and target feature distributions using adversarial training. However, in the setting with structured missingness that we consider, these methods can fail because the unobserved target subgroup may be incorrectly aligned to a visible source subgroup, a phenomenon termed as “collapse”. This occurs because adversarial alignment enforces marginal distribution matching without modeling hidden subpopulation structure, which can lead to biased predictions. Differently, our framework explicitly accounts for the unobservable source subpopulation, avoiding the collapse issue and providing more reliable adaptation in such scenarios.
@@ 3 Problem Setup and Notation
$R$ $Y$ AA $\bf X$} Sample Size Proportion Probability
Source 1 1 1 0 1 0 1 0 0 $n_{101}$ $n_{110}$ $n_{100}$ $p_{101}=\alpha_{01}\pi$ $p_{110}=\alpha_{10}\pi$ $p_{100}=\alpha_{00}\pi$ $\xi_{1}({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},A=1,R=1)\equiv 0$
@@
Note that this assumption is made without loss of generality, as alternative combinations, such as $(Y=0,A=1)$, $(Y=1,A=0)$, or $(Y=0,A=0)$, can be similarly assumed to have zero probability.
To characterize the distributional connection between the two domains, we impose a structured conditional invariance assumption:
@@
that is, the conditional distribution of features $\bf X$} given $(Y,A)$ remains the same across domains. This can be regarded as a conditional version, or, more nuanced version, of label shift where the marginal distribution of labels (now, the combination of both label and background) varies across domains (e.g., Du Plessis and Sugiyama, 2014; Garg et al., 2020; Iyer et al., 2014; Lipton et al., 2018; Nguyen et al., 2016; Tasche, 2017; Tian et al., 2023; Zhang et al., 2013; Lee et al., 2025a; b). This type of invariance assumptions was also postulated in other contexts such as in causal inference; see, e.g., Peters et al. (2016). It indicates, conditional on background $A$, the label shift assumption holds. It is equivalent to $p(R|{\bf X},Y,A)=p(R|Y,A)$, the independence between $R$ and $\bf X$}, conditional on $(Y,A)$. In practice, this assumption may be suitable in many applications. Below we give two examples to illustrate the rationality of this assumption. For instance, we aim to predict user clicks on advertisements for a new batch of users (target domain, $R=0$) using historical data (source domain, $R=1$). Conditional on the advertisement type $A$ and whether the user clicks $Y$, the distribution of browsing behavior features $\bf X$} is assumed to remain stable across time periods. This is because user clicks are fundamentally determined by ad content and user interests, not by the time period in which data are collected. As another example, suppose we have datasets from two hospitals ($R=1$ indicates the source hospital and $R=0$ indicates the target hospital). Here, $\bf X$} represents imaging features, $Y$ is the disease type, and $A$ denotes patient attributes such as gender or age group. Then, conditional on the disease type $Y$ and demographic attributes $A$, the distribution of imaging features $\bf X$} is expected to remain the same across hospitals. This is because imaging characteristics for a given disease and demographic group are not systematically altered by the hospital. The main difference between hospitals lies in sampling proportions rather than in conditional distributions.
@@ 4 Proposed Methodology
Our goal in this work is to correctly identify and successfully implement, under our UDA setting, the two background-specific predictive probabilities $\eta_{1}({\bf x})$ and $\eta_{0}({\bf x})$ and the overall predictive probability $\eta({\bf x})$, in the target domain. All of the three probabilities were precisely defined in Table 1.
4.1 The naive benchmarks
We consider two types of naive benchmarks. The first, termed Naive1 throughout, is to blindly apply the three source domain predictive probabilities $\xi_{1}({\bf x})$, $\xi_{0}({\bf x})$, and $\xi({\bf x})$ to the target, ignoring the structured missingness issue in this problem. More precisely, the first naive benchmark incorrectly treats $\xi_{1}({\bf x})$, $\xi_{0}({\bf x})$, and $\xi({\bf x})$ as $\eta_{1}({\bf x})$, $\eta_{0}({\bf x})$, and $\eta({\bf x})$, respectively. Note that, even though $\xi_{1}({\bf x})\equiv 0$ as indicated in Table 1, one can still use the observable data to implement the overall source domain predictive probability $\xi({\bf x})$ and the other background-specific predictive probability $\xi_{0}({\bf x})$ as
@@ 4.1 The naive benchmarks
The second type, termed Naive2, is to ignore the environment $A$ and simply impose a label shift assumption between the source and target domains. That is, one incorrectly specifies the following $\gamma({\bf x})$ as the overall target domain predictive probability $\eta({\bf x})$. After some simple algebra, one can compute
$\gamma({\bf x})=\frac{\frac{\beta_{11}+\beta_{10}}{\alpha_{11}+\alpha_{10}}\xi({\bf x})}{\frac{\beta_{11}+\beta_{10}}{\alpha_{11}+\alpha_{10}}\xi({\bf x})+\frac{\beta_{01}+\beta_{00}}{\alpha_{01}+\alpha_{00}}{\{1-\xi({\bf x})\}}}.$ (5)
@@ 4.2 Model adaptation from source to target
The most challenging aspect of this work is to adapt the model for the $A=1$ background since the component $(Y=1,A=1)$ is entirely absent in the source. Nevertheless, we can still correctly derive the three predictive probabilities for the target domain, as shown below.
@@
Define conditional probabilities $\tau_{0}({\bf x})=\hbox{pr}(A=1\mid{\bf X}={\bf x},R=0)$ and
@@
both of which can be implemented using the observed data in our UDA setting. Then the three predictive probabilities in the target domain are given by:
@@
The proof of this result is provided in Appendix A.1. Proposition 1 illustrated that, in general, the naive benchmarks presented in Section 4.1 fail. There are no explicit relations between $\eta_{1}({\bf x})$ and $\xi_{1}({\bf x})\equiv 0$ or between $\eta({\bf x})$ and $\xi({\bf x})$. For the relation between $\eta_{0}({\bf x})$ and $\xi_{0}({\bf x})$, they coincide only in the special case that $\beta_{10}/\alpha_{10}=\beta_{00}/\alpha_{00}$, which corresponds to a proportionality condition between the class-conditional densities across domains. Outside of this narrow scenario, the naive approach systematically misestimates the target posterior, leading to biased predictions.
This result also implies that model adaptation fundamentally relies on estimating the proportions of key subgroups in the target population. In particular, for individuals with $A=1$, one only needs to estimate $\beta_{01}$, while for those with $A=0$, it suffices to estimate the ratio $\beta_{10}/\beta_{00}$. Denote ${\bm{\beta}}=(\beta_{10},\beta_{00})^{\rm T}$. It can be seen that, accurate estimation of the parameter $\bm{\beta}$} in the target domain enables valid model adaptation across domains. Before developing methods for estimating $\bm{\beta}$} in Section 4.4, we first present some model identification considerations.